### v5

# Some new findings concerning value distribution of a pair of delay-differential polynomials

Abstract.
The paired Hayman’s conjecture of different types are considered. More accurately speaking, the zeros of a pair of $f^{n}L(z,g)-a_{1}(z)$ and $g^{m}L(z,f)-a_{2}(z)$ are characterized using different methods from those previously employed, where $f$ and $g$ are both transcendental entire functions, $L(z,f)$ and $L(z,g)$ are non-zero linear delay-differential polynomials, $\min\{n,m\}\geq 2$, $a_{1},a_{2}$ are non-zero small functions with relative to $f$ and $g$, or to $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$, respectively. These results give answers to three open questions raised by Gao, Liu[7] and Liu, Liu[15].
Key words and phrases:
Nevanlinna theory; Picard exceptional values; Entire functions; Paired Hayman’s conjecture; Delay-differential polynomials
2020 Mathematics Subject Classification: 39A05, 30D35
*Corresponding author.
1. Introduction
Let $f$ be a meromorphic function in the complex plane $\mathbb{C}$. Assume that the reader is familiar with the standard notation and basic results of Nevanlinna theory, such as $m(r,f),~{}N(r,f)$,${}T(r,f)$, see [8] for more details. A meromorphic function $g$ is said to be a small function of $f$ if $T(r,g)=S(r,f)$, where $S(r,f)$ denotes any quantity that satisfies $S(r,f)=o(T(r,f))$ as $r$ tends to infinity, outside of a possible exceptional set of finite linear measure. $\rho(f)=\underset{r\rightarrow\infty}{\lim\sup}\frac{\log^{+}T(r,f)}{\log r}$ and $\rho_{2}(f)=\underset{r\rightarrow\infty}{\lim\sup}\frac{\log^{+}\log^{+}T(r,f)}{\log r}$ are used to denote the order and the hyper-order of $f$, respectively. In general, the linear delay-differential polynomial of $f$ is defined by
$L(z,f)=\sum_{i=1}^{m}b_{i}(z)[f^{(\nu_{i})}(z+c_{i})],$
[…]
Definition.
Let $f$ be a meromorphic function. If $f-\alpha$ has finitely many zeros, then meromorphic function $\alpha$ is called a generalized Picard exceptional function of $f$. Furthermore, if $\alpha$ is also a small function of $f$, then we say $f$ has a generalized Picard exceptional small function $\alpha$. If $\alpha$ is a finite constant, then we say that $f$ has a generalized Picard exceptional value $\alpha$.
The well-known Picard theorem can be derived from the Nevanlinna’s second fundamental theorem. Here, we present the second fundamental theorem concerning three small functions[8, Theorem 2.5]: Let $f$ be a transcendental meromorphic function, and let $a_{j}$ $(j=1,2,3)$ be small functions of $f$, then
$T(r,f)<{\sum_{j=1}^{3}}\overline{N}(r,\frac{1}{f-a_{j}})+S(r,f).$
The generalizations of Nevanlinna’s second main theorem concerning $q~{}(\geq 3)$ small functions was given by Yamanoi[21], where $q$ is an integer. It is also easy to see a transcendental meromorphic function has at most two generalized Picard exceptional small functions.
The study of the generalized Picard exceptional values of complex differential polynomials has a long and rich history of research, including notable contributions such as Milloux’s inequality[8, Theorem 3.2], Hayman’s conjecture[9], Wiman’s conjecture[19]. In 1959, Hayman considered the value distribution of complex differential polynomials in his significant paper[9]. One of his results can be stated as follows.
Theorem A.
[9, Theorem 10] If $f$ is a transcendental entire function and $n~{}(\geq 2)$ is a positive integer, then $f^{n}f^{\prime}$ cannot have any non-zero generalized Picard exceptional values.
Clunie[5] proved that Theorem A is also true for the case $n=1$. The well-known Hayman’s conjecture is also presented in the same paper[9].
Hayman’s Conjecture.
[9] If $f$ is a transcendental meromorphic function and $n$ is a positive integer, then $f^{n}f^{\prime}$ cannot have any non-zero generalized Picard exceptional values.
Hayman’s conjecture has been solved completely. Hayman[9, Corollary to Theorem 9] obtained the proof for the case where $n\geq 3$. Mues[17] provided the proof for the case where $n=2$. Finally, using the theory of normal families, Bergweiler and Eremenko[2], Chen and Fang[4], and Zalcman[22] proved the case for $n=1$, respectively.
In 2007, Laine and Yang [10, Theorem 2] considered the generalized Picard exceptional values of complex difference polynomials, as outlined below, which can be viewed the difference analogues of Theorem A.
Theorem B.
[10] If $f$ is a transcendental entire function of finite order and $n~{}(\geq 2)$ is a positive integer, then $f^{n}f(z+c)$ cannot have any non-zero generalized Picard exceptional values, where $c$ is a non-zero constant.
For other results concerning the generalized Picard exceptional values of differences of meromorphic functions, please refer to [3, 6, 13, 14]. The generalized Picard exceptional small functions of delay-differential polynomials related to Hayman’s conjecture was considered by Liu, Liu and Zhou[16], and they obtained the following result.
Theorem C.
[16] Let $f$ be a transcendental meromorphic function of hyper-order $\rho_{2}(f)<1$, and let $c$ be a finite constant. If $n\geq 2k+6$ or if $n\geq 3$ and $f$ is a transcendental entire function, then $f^{n}f^{(k)}(z+c)$ can not have any non-zero generalized Picard exceptional small function $a$ with respect to $f$.
Later, the condition $n\geq 2k+6$ in Theorem C has been weakened by Laine et al.[11] to $n>k+4$.
2. Hayman’s conjecture on a pair of delay-differential polynomials
In 2022, Gao and Liu[7] considered the Hayman’s conjecture regarding a pair of delay-differential polynomials.
Theorem D .
[7] Let $f$ and $g$ be both transcendental meromorphic functions, $n$ and $k$ be non-negative integers, $\alpha$ be a non-zero small function with respect to $f$ and $g$. If one of the following conditions is satisfied:
  • (1)
    $f$ and $g$ are both entire functions, $n\geq 3$, $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,h)=h^{(k)}(z+c)$ or $L(z,h)=h(z+c)-h(z)$;
  • (2)
    $f$ and $g$ are meromorphic functions, $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,h)=h(z+c)$, $n\geq 4$ or $L(z,h)=h(z+c)-h(z)$, $n\geq 5$ or $L(z,h)=h^{(k)}(z+c)$, $n\geq k+4$, respectively;
  • (3)
    $f$ and $g$ are entire functions, $n\geq 3$, $L(z,h)=h^{(k)}(z)$, or $f$ and $g$ are meromorphic functions, $n\geq k+4$, $L(z,h)=h^{(k)}(z)$, respectively;
then $f^{n}L(z,g)$ and $g^{n}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$.
Obviously, if $f\equiv g$, then Theorem D may reduce to Hayman’s conjecture of different types. Regarding Theorem D, Gao and Liu posed the following question for further investigation.
Question 1.
[7, Question 1] Can we reduce $n\geq 3$ to $n\geq 2$ in Theorem D for entire functions $f$ and $g$? And what is the sharp value $n$ for meromorphic functions $f,g$?
By considering a broader class of a pair of delay-differential polynomials, we give a positive answer to the first part of Question 1.
Theorem 2.1.
Let $f$ and $g$ be both transcendental entire functions, $\min\{n,m\}\geq 2$, $\alpha$ be a non-zero small function with respect to $f$ and $g$. Then, $f^{n}L(z,g)$ and $g^{m}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small entire functions of $f$ and $g$ as its coefficients.
Remark 1.
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are non-zero linear delay-differential polynomial with small entire functions of $f$ and $g$ as their coefficients, $\alpha$ is a non-zero small funtion with respect to $f$ and $g$, then $f^{n}L(z,g)$ and $g^{n}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$ by using the similar proof of Theorem 2.1. Combining this result with Theorem 2.1, we give a positive answer to Question 1 for entire functions $f$ and $g$.
Remark 2.
Obviously, if $f\equiv g$ is an entire function, Theorem 2.1 can be reduced to Theorems A, B and [12, Corollary 1.7], and improves Theorems C. What’s more, our approaches differ significantly from the methods outlined in Theorems A, B, C and D.
Remark 3.
The restriction that non-zero Picard exceptional small functions of $f$ and $g$ can not be removed. For example, let $f(z)=e^{z}$, $g(z)=e^{2z},$ $L(z,f)=f^{\prime\prime}+f^{\prime}$, $L(z,g)=g^{\prime\prime}+g^{\prime}$, then $f^{3}L(z,g)=6e^{5z}$, $g^{4}L(z,f)=2e^{9z}$ and 00 is a Picard exceptional value of $f^{3}L(z,g)$ and $g^{5}L(z,f)$.
Remark 4.
If $m=n=1$, then Theorem 2.1 does not holds. For example, let $f(z)=e^{-p(z)}$, $g(z)=e^{p(z)}$, where $p$ is a non-constant polynomial, then $fg-2=-1$ has no zeros.
From Theorem D, it is evident that a small function $\alpha$ respect to $f$ and $g$ can not serve as a common generalized Picard exceptional function for both $f^{n}L(,g)$ and $g^{n}L(z,f)$. Naturally, this prompts the question[15, Question 1.1]: Given any two transcendental meromorphic functions $F(z)$ and $G(z)$, what can we obtain for their common or different generalized Picard exceptional values or small functions of $F$ and $G$? The direct consideration of this question is rather difficult, thus Liu and Liu[15] raised the following question.
Question 2.
[15, Question 2.1] Let $f$ and $g$ be two transcendental meromorphic functions and $m$, $n$ be positive integers. Do $f^{n}g$ and $g^{m}f$ have simultaneously the generalized Picard exceptional values or small functions?
Liu and Liu[15] considered Question 2 and obtained partial results.
Theorem E .
[15, Theorem 2.1 and Theorem 3.1] Let f and g be transcendental entire functions.
  • (1)
    If $\min\{m,n\}\geq 2$, then $f^{n}g$ and $g^{m}f$ can not have simultaneously non-zero generalized Picard exceptional value.
  • (2)
    If $n\geq 2$ then $f^{n}g$ and $gf$ can not have simultaneously non-zero generalized Picard exceptional value except that $f(z)=se^{T}(z)$ and $g(z)=s^{-2}a_{1}e^{-2T(z)}+a_{2}s^{-1}e^{-T(z)}$, where $T(z)$ is an entire function and $s$ is a complex constant. In this case, $a_{1}$ is the Picard exceptional value of $f^{2}g$ and $a_{2}$ is the Picard exceptional value of $fg$.
  • (3)
    If $\min\{m,n\}\geq 3$, $k$ is a positive integer, then $f^{n}g^{(k)}$ and $g^{m}f^{(k)}$ can not have simultaneously non-zero Picard generalized exceptional small functions.
Liu and Liu asked the following question related Theorem E-$(3)$.
Question 3.
[15, Remark 3.2] Can we reduce $n\geq 3$ to $n\geq 2$ in Theorem E-$(3)$?
We consider Questions 2 and 3 and obtain the following results.
Theorem 2.2.
Let $f$ and $g$ be both transcendental entire functions.
  • (1)
    If $\min\{m,n\}\geq 2$, then $f^{n}L(z,g)$ and $g^{m}(z)L(z,f)$ cannot have simultaneously non-zero generalized Picard exceptional small functions, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small entire functions of $f$ and $g$ as its coefficients.
  • (2)
    If $n\geq 2$, then $f^{n}g$ and $gf$ can not have simultaneously non-zero generalized Picard exceptional value except that $f(z)=A(z)e^{B(z)}$, $g(z)=A_{4}(z)e^{-nB(z)}+A_{3}(z)e^{-B(z)}$, where $A,A_{3},A_{4}$ are small entire function of $f$ and $g$, $B$ is an entire function. What’s more, $A^{n}A_{4}$ is the Picard exceptional small function of $f^{n}g$, $AA_{3}$ is the Picard exceptional small function of $fg$.
Remark 5.
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are linear delay-differential polynomial with small entire functions of $f$ and $g$ as its coefficients, then Theorem 2.2-$(1)$ still holds by using the similar proof of Theorem 2.2.
Remark 6.
If $L(z,f)=f^{(k)}$ and $L(z,g)=g^{(k)}$, then Theorem 2.2-(1) gives a positive answer to Question 3.
Remark 7.
Note that $f(z)=e^{z}$ and $g(z)=2e^{-3z}+e^{-z}$ satisfy the relationships $f^{3}(z)g(z)=2+e^{2z}$ and $g(z)f(z)=2e^{-2z}+1$. Consequently, $2$ is a Picard exceptional value of $f^{3}g$, $1$ is a Picard exceptional value of $gf$. However, $g(z)=2e^{-3z}+e^{-z}$ does not conform to the form $g(z)=s^{-2}a_{1}e^{-2T(z)}+a_{2}s^{-1}e^{-T(z)}$ specified in Theorem E-$(2)$. This suggests that there may be a gap in the proof of Theorem E-$(2)$. Our Theorem 2.2-$(2)$ rectify the gap using different methods from those in Theorem E.
Remark 8.
An example of Theorem 2.2-$(2)$ in the context of small functions is given as follows: $f(z)=e^{z^{2}+z}e^{z^{5}}$, $g(z)=e^{-5z^{5}}+e^{z^{3}}e^{-z^{5}}$, which satisfies $f^{5}g=e^{5z^{2}+5z}+e^{4z^{5}}e^{5z^{2}+5z+z^{3}}$, $fg=e^{z^{2}+z}e^{-4z^{5}}+e^{z^{2}+z+z^{3}}$. Here, $e^{5z^{2}+5z}$ is a Picard exceptional small function of $f^{5}g$, $e^{z^{2}+z+z^{3}}$ is a Picard exceptional small function of $fg$.
Remark 4 can also be used to demonstrate that Theorem 2.2 does not hold when $n=m=1$. Combining this with Theorem 2.2, we obtain the following corollary, which completely solves Question 2 for the case where $f$ and $g$ are entire functions.
Corollary 2.3.
Let $f$ and $g$ be two transcendental entire functions and $m$, $n$ be positive integers, then $f^{n}g$ and $g^{m}f$ can not have simultaneously the generalized Picard exceptional small functions except $\min\{m,n\}=1$.
3. Proof of Theorems 2.1 and 2.2
Proof of Theorem 2.1.
Without loss of generality, we will prove Theorem 2.1 in two cases: whether $g$ is a small function of $f$ or not.
Case 1.
$g$ is a small function of $f$. Since $\alpha$ is a small function of $f$, then $\alpha$ is also a small function of $f^{n}L(z,g)$. We claim $f^{n}L(z,g)-\alpha$ has infinitely many zeros. Otherwise, if $f^{n}L(z,g)-\alpha$ has finitely many zeros, by the Weierstrass’s factorization theorem[18, pp. 145], we have
(3.1) $f^{n}(z)L(z,g)-\alpha(z)=p(z)e^{b(z)},$
where $p$ is the infinite products formed by the zeros and poles of $f^{n}L(z,g)-\alpha$, $b$ is an entire function, and they satisfy $T(r,p)=S(r,f^{n}L(z,g))=S(r,f)$ and $T(r,e^{b})=T(r,f^{n}L(z,g)+S(r,f)=O(T(r,f))+S(r,f)$.
Differentiating the equation (3.1), we get
(3.2) $nf^{n-1}f^{\prime}L(z,g)+f^{n}L^{\prime}(z,g)-\alpha^{\prime}=p_{1}e^{b},$
where $p_{1}=p^{\prime}+b^{\prime}p$. We claim $p^{\prime}+b^{\prime}p\not\equiv 0$. Otherwise $p^{\prime}+b^{\prime}p\equiv 0$, by integrating, we can obtain $p(z)=c_{1}e^{-b(z)}$, where $c_{1}$ is a non-constant. Which is impossible, since $T(r,p)=S(r,f)$.
By eliminating $e^{b}$ from equations (3.1) and (3.2), we can obtain
(3.3) $p_{1}f^{n}L(z,g)-pnf^{n-1}f^{\prime}L(z,g)-pf^{n}L^{\prime}(z,g)=p_{1}\alpha-p\alpha^{\prime}.$
We claim $p_{1}\alpha-p\alpha^{\prime}\not\equiv 0$. Otherwise $p_{1}\alpha-p\alpha^{\prime}\equiv 0$, by integrating, we can obtain $\frac{\alpha}{p}=c_{2}e^{b}$, which is impossible, since $T(r,p)=S(r,f)$ and $T(r,\alpha)=S(r,f)$.
Suppose $N(r,\frac{1}{f})\not=S(r,f)$, since coefficients of the equation (3.3) are small functions of $f$, then there exists a zero $z_{0}$ of $f$ such that the coefficients of the equation (3.3) are neither zero nor infinite at that point. Substituting $z_{0}$ into the equation (3.3), we easily obtain a contradiction. Therefore $N(r,\frac{1}{f})=S(r,f)$. By the Weierstrass’s factorization theorem, let $f(z)=A(z)e^{B(z)}$, where $A$ and $B$ are entire function, such that $T(r,A)=S(r,f)$. By the second fundamental theorem concerning three small functions and (3.1), we have
$T(r,f^{n})<\overline{N}(r,\frac{1}{f^{n}})+\overline{N}(r,\frac{1}{f^{n}-\frac{\alpha}{L(z,g)}})+\overline{N}(r,f^{n})+S(r,f)=S(r,f),$
which is impossible. Therefore $f^{n}L(z,g)-\alpha$ has infinitely many zeros. This means that $\alpha$ cannot be a common non-zero generalized Picard exceptional small function for both $f^{n}L(z,g)$ and $g^{m}L(z,f)$.
Case 2.
$g$ is not a small function of $f$. We don’t need to consider the case $f$ is a small function of $g$, since this case is as the same as the Case 1. Therefore, in this case $T(r,f)=O(T(r,g))$, which means $S(r,f)=S(r,g)$. Let $f^{n}L(z,g)=f_{1}$, $g^{m}L(z,f)=f_{2}$.
By the logarithmic derivative lemma[8, Theorem 2.2], it is easily to get
$T(r,f^{n})$ $=m(r,f^{n})=m(r\frac{f^{n}}{L(z,g)}L(z,g))\leq m(r,f_{1})+m(r,\frac{L(z,g)}{g})+m(r,g)+S(r,g)$
$\leq m(r,f_{1})+m(r,g)+S(r,g),$
(3.4) $T(r,f_{1})$ $\leq nm(r,f)+m(r,g)+S(r,g).$
$T(r,g^{m})$ $=m(r,g^{m})=m(r\frac{g^{m}}{L(z,f)}L(z,f))\leq m(r,f_{2})+m(r,\frac{L(z,f)}{f})+m(r,f)+S(r,f)$
$\leq m(r,f_{2})+m(r,f)+S(r,f),$
(3.5) $T(r,f_{2})$ $\leq mm(r,g)+m(r,f)+S(r,f).$
We only need to consider $m(r,g)\leq m(r,f)$, since the case $m(r,f)\leq m(r,g)$ can be studied by the same methods.
Since $m(r,g)\leq m(r,f)$, then from (2), we get $(n+1)m(r,f)+S(r,f)\geq m(r,f_{1})\geq(n-1)m(r,f)$, which means $\alpha$ is a small function of $f_{1}$.
Below, we will consider two subcases: $m(r,g)\leq m(r,f)$ and whether $f_{2}$ is a small function of $f$ or not.
Subcase 2.1.
$m(r,g)\leq m(r,f)$ and $f_{2}$ is not a small function of $f$. Since $f_{2}$ is not a small function of $f$, then $T(r,f_{2})=O(T(r,f))$, $\alpha$ is a small function of $f_{2}$. We claim at least one of $f^{n}L(z,g)-\alpha$ and $g^{m}L(z,f)-\alpha$ has infinitely many zeros. Otherwise, suppose $f^{n}L(z,g)-\alpha$ and $g^{m}L(z,f)-\alpha$ have finitely many zeros. Since $\alpha$ is a small function of $f_{1}$ and $f_{2}$, by the Weierstrass’s factorization theorem, we have
(3.6) $f^{n}L(z,g)-\alpha=p_{1}e^{b_{1}(z)},$
where $p_{1}$ is a small functions of $f^{n}L(z,g)$, $b_{1}(z)$ is an entire function and
(3.7) $g^{m}L(z,f)-\alpha=p_{2}e^{b_{2}(z)},$
where $p_{2}$ is a small functions of $g^{m}L(z,f)$, $b_{2}(z)$ is an entire function.
From (2) and (2), we see the small functions of $f_{1}$ and $f_{2}$ are also the the small functions of $f$ and $g$, therefore $p_{1}$ and $p_{2}$ are small functions of $f$ and $g$. Using the same methods of Case 1 to (3.6) and (3.7), we also can get $N(r,\frac{1}{f})=S(r,f)$ and $N(r,\frac{1}{g})=S(r,g)$. By the Weierstrass’s factorization theorem, we can get $f(z)=A_{1}(z)e^{B_{1}(z)}$, where $A_{1}$ and $B_{1}$ are entire function, such that $T(r,A_{1})=S(r,f)$, $g(z)=A_{2}(z)e^{B_{2}(z)}$, where $A_{2}$ and $B_{2}$ are entire function, such that $T(r,A_{2})=S(r,f)$. Let $\frac{L(z,f)}{f(z)}=a_{1}$, By applying the logarithmic derivative lemma, we can obtain $m(r,\frac{L(z,f)}{f})\leq S(r,f)$. Since $N(r,\frac{1}{f})=S(r,f)$, we have $T(r,\frac{L(z,f)}{f})=m(r,\frac{L(z,f)}{f})+N(r,\frac{L(z,f)}{f})=S(r,f)$. Thus $a_{1}$ is a small function of $f$. Using the same derivation, we get $L(z,g)=g(z)a_{2}$, where $a_{2}$ is a small function of $g$. Substituting the expressions for $f$, $L(z,f)$, $g$ and $L(z,g)$ into the equations (3.6) and (3.7), we obtain
(3.8) $\begin{cases}a_{3}e^{nB_{1}+B_{2}}-\alpha=p_{1}e^{b_{1}},\\ a_{4}e^{mB_{1}+B_{2}}-\alpha=p_{2}e^{b_{2}},\end{cases}$
where $a_{3}$ and $a_{4}$ are small function of $f$ and $g$. By the second fundamental theorem concerning three small functions, we have
$T(r,a_{3}e^{nB_{1}+B_{2}})$ $<\overline{N}(r,\frac{1}{a_{3}e^{nB_{1}+B_{2}}})+\overline{N}(r,\frac{1}{a_{3}e^{nB_{1}+B_{2}}-\alpha})+\overline{N}(r,{a_{3}e^{nB_{1}+B_{2}}}))+S(r,f)$
$=S(r,f),$
which is impossible. Therefore at least one of $f^{n}L(z,g)-\alpha$ and $g^{n}L(z,f)-\alpha$ has infinitely many zeros. This means that $\alpha$ cannot be a common non-zero generalized Picard exceptional small function for both $f^{n}L(z,g)$ and $g^{m}L(z,f)$.
Subcase 2.2.
$m(r,g)\leq m(r,f)$ and $f_{2}$ is a small function of $f$. We claim $f^{n}L(z,g)-\alpha$ has infinitely many zeros. Otherwise, $f^{n}L(z,g)-\alpha$ has finitely many zeros. As the same as the Case 1, we get equations (3.1)-(2) and $f(z)=A_{3}(z)e^{B_{3}(z)}$, where $A_{3}$ and $B_{3}$ are entire function, such that $T(r,A_{3})=S(r,f)$.
By applying the logarithmic derivative lemma, we can obtain $m(r,\frac{L(z,f)}{f})\leq S(r,f)$. Since $N(r,\frac{1}{f})=S(r,f)$, we have $T(r,\frac{L(z,f)}{f})=m(r,\frac{L(z,f)}{f})+N(r,\frac{L(z,f)}{f})=S(r,f)$. Thus $\frac{L(z,f)}{f}=a_{5}$, where $a_{5}$ is a small function of $f$. Therefore $g^{m}=\frac{f_{2}}{L(z,f)}=f_{2}\frac{1}{a_{5}f}$, which means $g=A_{4}e^{\frac{-B_{3}}{m}}$, which $A_{4}$ is a small function of $f$. As the same as $L(z,f)$, we have $L(z,g)=a_{6}g$, where $a_{6}$ is a small function of $g$.
Substituting the expressions for $f$, $g$ and $L(z,g)$ into the equation (3.1), we obtain
$a_{7}e^{\frac{n-1}{m}B_{3}}-\alpha=pe^{b},$
where $a_{7}$ is a small function of $f$. Since $n\geq 2$, we have $T(r,a_{7}e^{\frac{n-1}{m}B_{3}})=O(T(r,f))+S(r,f)$. By the second fundamental theorem concerning three small functions, we have
$T(r,a_{7}e^{\frac{n-1}{m}B_{3}})<\overline{N}(r,\frac{1}{a_{7}e^{\frac{n-1}{m}B_{3}}})+\overline{N}(r,\frac{1}{a_{7}e^{\frac{n-1}{m}B_{3}}-\alpha})+\overline{N}(r,{a_{7}e^{\frac{n-1}{m}B_{3}}}))+S(r,f)=S(r,f),$
which is impossible. Therefore $f^{n}L(z,g)-\alpha$ has infinitely many zeros. This means that $\alpha$ cannot be a common non-zero generalized Picard exceptional small function for both $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$.
Proof of Theorem 2.2.
We will prove the Theorem 2.2 by using similar ideas in proving Theorem 2.1, we divide the proof into two cases.
Case 1.
Suppose $\min\{m,n\}\geq 2$ and $f^{n}L(z,g)$ and $g^{m}L(z,f)$ have simultaneously non-zero generalized Picard exceptional small functions. By Weierstrass’s factorization theorem, we assume that
(3.9) $\begin{cases}f^{n}L(z,g)-a_{1}(z)=p_{1}(z)e^{b_{1}(z)},\\ g^{m}L(z,f)-a_{2}(z)=p_{2}(z)e^{b_{2}(z)},\end{cases}$
where $a_{1}$ and $b_{1}$ are non-zero small functions of $f^{n}L(z,g)$, $a_{2}$ and $b_{2}$ are non-zero small functions of $g^{m}L(z,f)$.
If $g$ is a small function of $f$. Since $a_{1}$, $p_{1}$ are small functions of $f^{n}L(z,g)$, then $a_{1}$ and $p_{1}$ are also small functions of $f$. If $g$ is not a small function of $f$. Since $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f^{n}L(z,g)$ and $g^{m}L(z,f)$, receptively, it is easy to get $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f$ and $g$, receptively. These two situations are similar to case 1 and case 2 of Theorem 2.1, then we can also derive a contradiction through the same method. Therefore we omit the proof.
Case 2.
Suppose $f^{n}g$ and $gf$ have simultaneously non-zero generalized Picard exceptional small functions. By Weierstrass’s factorization theorem, we assume that
(3.10) $\begin{cases}f^{n}g-a_{1}(z)=p_{1}(z)e^{b_{1}(z)},\\ gf-a_{2}(z)=p_{2}(z)e^{b_{2}(z)},\end{cases}$
where $a_{1}$ and $b_{1}$ are non-zero small functions of $f^{n}g$, $a_{2}$ and $b_{2}$ are non-zero small functions of $gf$.
If $g$ is a small function of $f$, we can derive a contradiction through the same method of Case 1. Therefore, we only need to consider $g$ is not a small function of $f$. Thus $T(r,f)=O(T(r,g))$, $S(r,f)=S(r,g)$. Since $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f^{n}L(z,g)$ and $g^{m}L(z,f)$, receptively, it is easy to get $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f$ and $g$.
Since $a_{1}$ is a small function of $f^{n}L(z,g)$, then $a_{1}$ is also a small function of $f$. Using the same methods of Case 1 of the proof of Theorem 2.1 to the first equation (3.10), we also get $N(r,\frac{1}{f})=S(r,f)$. By the Weierstrass’s factorization theorem, we can get $f(z)=A(z)e^{B(z)}$, where $A$ is a small function of $f$, $B$ is an entire function. Substituting the expression for $f$ into the second equation of (3.10), we get
(3.11) $g(z)=A_{2}(z)e^{b_{2}(z)-B(z)}+A_{3}(z)e^{-B(z)},$
where $A_{2}$, $A_{3}$ are small functions of $f$ and $g$. Substituting the expressions for $f$ and $g$ into the first equation of (3.10), we get
(3.12) $a_{3}e^{(n-1)B+b_{2}}+a_{4}e^{(n-1)B}-a_{1}=p_{1}e^{b_{1}},$
where $a_{3},a_{4}$ are non-zero small functions of $f$ and $g$.
Since $a_{1},a_{3},a_{4}$ are small functions of $f$, then they are also the small functions of $e^{b_{1}}$ and $e^{(n-1)B}$. Therefor, by [20, Theorem 1.51] and (3.12), we get $a_{3}e^{(n-1)B+b_{2}}=a_{1}$, which means $T(r,e^{(n-1)B+b_{2}})=S(r,f)$. Substituting the expressions for $e^{b_{2}}$ into (3.11), we get $g(z)=A_{4}(z)e^{-nB(z)}+A_{3}(z)e^{-B(z)}$, where $A_{4}$ is a non-zero small function of $g$. Substituting $f$ and $g$ in (3.10) and by [20, Theorem 1.51], we get $A^{n}A_{4}=a_{1}$, $AA_{3}=a_{2}$.